OpenStudy (anonymous):

Suppose y varies directly as x, and y = 10 when x = 6. What is the value of x when y = 20?

OpenStudy (anonymous):

direct variation is y=kx so we have a y and x. we need to find k so substitute y and x 10=6k k=10/6 or 5/3 now we substitute k into the equation y=5/3x we can now substitue y 20=5/3x multiply by 3/5 or divide by 5/3 x=12